Permutations are just an application of fundamental counting principle. Combinations are just a application of Permutations and the counting principle. So would you say that Combinations are also just an application of the counting principle?
Answer by stanbon(75887)

Yes and no.
8C3 = [8*7*6]/[3*2*1]=(8/3)(7/2)(6/1) = 56
It looks like a counting principal application but
you would have a hard time putting those fractions into words
that describe the number of ways to select a group of three from
a group of eight
Cheers,
Stan H